# Neural Geometry Processing via Spherical Neural Surfaces

### v1

@@ 3.2. Computing Differential Quantities using SNS
(5) $R=\begin{pmatrix}\cos\theta\cos\phi&-\sin\phi\\ \cos\theta\sin\phi&\cos\phi\\ -\sin\theta&0\\ \end{pmatrix}.$
This matrix is an orthonormal version of the usual spherical polar coordinates Jacobian. We avoid degeneracy at the poles by selectively re-labeling the $x$, $y$, and $z$ coordinates so that $\theta$ is never close to zero or $\pi$.
Now, from $\mathbf{J}^{\textit{local}}_{S}$, we can simply write the First Fundamental Form (FFF) of the sphere-to-surface map as:
(6) $\mathbf{I}_{S}=\begin{pmatrix}E&F\\ F&G\\ \end{pmatrix}=(\mathbf{J}^{\textit{local}}_{S})^{T}\,\mathbf{J}^{\textit{local}}_{S}.$
To compute the Second Fundamental Form (SFF), we first compute the second partial derivatives of the parametrization with respect to the local co-ordinates, and then find the dot products of the second derivatives with the normal, $\mathbf{n}=\mathbf{n}_{S_{\theta}}$. The second derivatives are:
(7) $S_{uu}=\mathbf{J}(S_{u})S_{u},\quad S_{uv}=\mathbf{J}(S_{u})S_{v},\quad S_{uu}=\mathbf{J}(S_{v})S_{v}.$
We can now express the SFF as:
(8) $\mathbf{II}_{S}=\begin{pmatrix}e&f\\ f&g\\ \end{pmatrix}=\begin{pmatrix}S_{uu}\cdot\mathbf{n}&S_{uv}\cdot\mathbf{n}\\ S_{uv}\cdot\mathbf{n}&S_{vv}\cdot\mathbf{n}\\ \end{pmatrix}.$
In terms of the FFF and SFF, the Gauss curvature is
(9) $K=\frac{eg-f^{2}}{EG-F^{2}},$
@@
(10) $H=\frac{Eg-2Ff+Ge}{2(EG-F^{2})}.$
Furthermore, we can algebraically solve the quadratic equation $\text{det}(\mathbf{I}_{S}-\lambda\mathbf{II}_{S})=0$, to obtain explicit expressions for the principal curvatures, that we then use to find the principal curvature directions directly on the Spherical Neural Surface. In Figure 3, we compare our differential estimates against analytical computations for an analytic surface; in Figure 6 we also show curvature estimates and principal curvature directions on other SNS approximations of input triangle meshes (ground truth values are not available).
Figure 3. First and Second Fundamental Forms. Here, we display the differential quantities computed using our framework, on an SNS that has been overfitted to the mesh of a surface with an analytical parametrization (bottom) and we show that the results are extremely close to the differential quantities that have been computed symbolically, using Matlab.
4. Mathematical Background
@@
4.1. The Laplace Beltrami Operator on Smooth Surfaces
The (Euclidean) Laplacian $\Delta$ of a scalar field $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is defined as the divergence of the gradient of the scalar field. Equivalently, $\Delta f$ is the sum of the unmixed second partial derivatives of the field $f$:
(11) $\Delta f:=\nabla\cdot\nabla f=\sum_{i=1}^{n}\frac{\partial^{2}f}{\partial^{2}x_{i}}.$
@@ 4.1. The Laplace Beltrami Operator on Smooth Surfaces
(12) $\Delta_{\Sigma}f:=\nabla_{\Sigma}\cdot\nabla_{\Sigma}f.$
The surface gradient of the scalar field $f$ (written $\nabla_{\Sigma}f$), and the surface divergence of a vector field ${\mathbf{F}}:\Sigma\rightarrow\mathbb{R}^{3}$ (written $\nabla_{\Sigma}\cdot F$), can be computed by smoothly extending the fields to fields $\tilde{f}$ and $\tilde{\mathbf{F}}$ respectively, which are defined on a neighborhood of $\Sigma$ in $\mathbb{R}^{3}$.
The surface gradient is defined as the projection of the gradient of $\tilde{f}$ into the local tangent plane:
(13) $\nabla_{\Sigma}f=\nabla\tilde{f}-(\nabla\tilde{f}\cdot\mathbf{n})\mathbf{n},$
where $\mathbf{n}$ is the unit surface normal at the point. The surface divergence of the vector field $\mathbf{F}$ is defined as
(14) $\nabla_{\Sigma}\cdot\mathbf{F}=\nabla\cdot\tilde{\mathbf{F}}-\mathbf{n}^{T}{\mathbf{J}}(\tilde{\mathbf{F}})\mathbf{n},$
@@
By expanding out the definitions of surface gradient and surface divergence, we can derive an alternative formula for the surface Laplacian ($\Delta_{\Sigma}f$) in terms of the (Euclidean) Laplacian ($\Delta\tilde{f}$), the gradient ($\nabla\tilde{f}$) and the Hessian ($\mathbf{H}(\tilde{f})=\mathbf{J}(\nabla\tilde{f})^{T}$) as,
(15) $\Delta_{\Sigma}f=\Delta\tilde{f}-2H\nabla\tilde{f}\cdot\mathbf{n}-\mathbf{n}^{T}\mathbf{H}(\tilde{f})\mathbf{n}.$
The dependence on the surface is captured by the normal function $\mathbf{n}$, and the mean curvature $H$. Please refer to (Reilly, 1982; Xu and Zhao, 2023) for a derivation. In the case when $\tilde{f}$ is a ‘normal extension’ — i.e., it is locally constant in the normal directions close to the surface — then the second and third terms disappear, meaning that $\Delta_{\Sigma}f$ is consistent with $\Delta\tilde{f}$.
If we substitute in the coordinate function $\mathbf{x}$, then the first and third terms disappear in the above equation, and we are left with the familiar equation,
(16) $\Delta_{\Sigma}\mathbf{x}=-2H\mathbf{n}.$
@@
4.2. Spectrum of the Laplace Beltrami Operator
@@
In the discrete mesh case, the eigenfunctions of the Laplace-Beltrami operator are computed by diagonalizing a sparse matrix (Lévy and Zhang, 2010; Bunge et al., 2023) based on the input mesh. In the continuous case, however, we no longer have a matrix representation for $\Delta_{\Sigma}$. Instead, we exploit a functional analysis result that describes the first $k$ eigenfunctions of $\Delta_{\Sigma}$ as the solution to a minimization problem, as described next.
@@
(19) $Q_{\Sigma}(f):=\frac{\left\|\nabla_{\Sigma}f\right\|_{L^{2}(\Sigma)}^{2}}{\left\|f\right\|^{2}_{L^{2}(\Sigma)}}.$
@@
(20) $\textit{such that}\quad\left<\Psi_{i},\Psi_{j}\right>_{L^{2}(\Sigma)}=0\quad\text{for all}\quad i\neq j.$
@@
Figure 4. Neural Spectral Basis against Analytical Solutions. The sphere is a genus-0 surface for which the eigenspaces of the Laplace-Beltrami operator are very well understood: the eigenspaces are the spaces spanned by linear combinations of the spherical harmonics for each frequency. This means that each of our optimized eigenfunctions should belong to one of these eigenspaces, and therefore, the Rayleigh quotient should match the energy level of the corresponding spherical harmonics. For evaluation, we compute the Rayleigh quotient using a large number of samples for Monte Carlo integration, since we cannot analytically compute the integral. For reference, we present the ground truth Rayleigh quotients, in green.
@@
(21) $\Delta_{\Sigma}\Psi_{i}(x)=-Q_{\Sigma}(\Psi_{i})\Psi_{i}(x)\quad\forall\,x\in\Sigma.$
@@
Physically, the eigenfunctions of $\Delta_{\Sigma}$ represent the fundamental modes of vibration of the surface $\Sigma$. The Rayleigh Quotient of $\Psi_{0}$ is zero, and as the Rayleigh quotient increases, the vibrational modes increase in energy and the eigenfunctions produce more ‘intricate’ (high frequency) patterns.
Figure 5. Encoding Scalar Fields on Neural Surfaces. We encode scalar fields on the sphere (and corresponding extensions into $\mathbb{R}^{3}$) as MLPs
@@
$g_{\eta}:\mathbb{R}^{3}\rightarrow\mathbb{R}$. Because $S_{\theta}$, considered as a parametrization, is bijective, then we can implicitly define any smooth scalar field $h$ on the surface in terms of one of the functions $g_{\eta}$ (restricted to $\mathbb{S}^{2}$). We visualize the one-to-one correspondence between scalar fields defined on the surface, and scalar fields defined on the sphere, by using vertex-colors to display the fields.
@@
Figure 6. First and Second Fundamental Forms. An SNS represents a smooth, seamless parametrization of a genus-0 surface, and we can compute the partial derivatives of this parametrization using automatic differentiation. We compute the normals and the First Fundamental Form of the parametrization from the first derivatives. Subsequently, computing the second derivatives enables us to construct the Second Fundamental Form of the parametrization. Hence, we compute Mean Curvature, Gaussian Curvature, and Principal Curvature Directions at any point without any unnecessary discretization error.
5. Spectral Analysis using SNS
@@
To find the continuous eigenmodes of the Laplace-Beltrami operator on an SNS, we require a continuous and differentiable representation for scalar functions defined on the surface. In our framework, we represent such smooth scalar fields on $\mathbb{S}^{2}$ by MLPs, $g_{\eta}$, from $\mathbb{R}^{3}$ to $\mathbb{R}$. We only ever evaluate the MLP $g_{\eta}$ on $\mathbb{S}^{2}\in\mathbb{R}^{3}$, but because the domain is $\mathbb{R}^{3}$ then it defines an extension of the scalar field, and this allows us to compute $\nabla g_{\eta}$. Then, if $S_{\theta}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ is an SNS, any smooth scalar field $h$ defined on $S_{\theta}$ can be defined implicitly by the equation,
(22) $h\circ S_{\theta}=g_{\eta}.$
Applying the chain rule to the implicit equation, we can compute the gradient of $h$:
(23) $\nabla h=(\mathbf{J}(S_{\theta})^{T})^{-1}\nabla g_{\eta},$
@@
To optimize for a (smooth) scalar function $g_{\eta_{k}}$ so that $h$ approximates the $k$th non-constant eigenfunction of the Laplace-Beltrami operator, we use gradient descent to optimize the weights, $\eta_{k}$. We use a combination of two loss terms:
(24) $\mathcal{L}_{Rayleigh}:=Q_{S}(h)\quad\text{and}\quad\mathcal{L}_{ortho}:=\sum_{i=0}^{k-1}\left<h,h_{i}\right>_{L^{2}(S)}^{2},$
where we sequentially compute the eigenfunctions, and $h_{i}$ is the $i$th smallest eigenfunction to be found. Recall that we represent each such eigenfunction using a dedicated MLP (except for $h_{0}$, which is the constant eigenfunction). The loss term $\mathcal{L}_{ortho}$ encourages the scalar field $h$ to be orthogonal to all previously-found eigenfunctions (including the constant eigenfunction). The quantity $Q_{S_{\theta}}(h)$ depends on the Jacobian of $S_{\theta}$, through $\nabla_{S_{\theta}}h$ and $\nabla h$. However, since the surface does not change during optimization and $\theta$ stays fixed, we precompute $(\mathbf{J}(S_{\theta})^{T})^{-1}$. Note that, if we were to optimize over the subspace of the unit-norm functions in $L^{2}(\Sigma)$, then we could replace the Rayleigh Quotient with the Dirichlet energy. However, it was a more natural design choice for us to parametrize the entire space $L^{2}(\Sigma)$ and optimize the Rayleigh Quotient, rather than to parametrize only the space of unit-norm functions in $L^{2}(\Sigma)$ and minimize the Dirichlet Energy.
Additionally, for stability, we included a regularization term designed to keep the $L^{2}$-norm of $h$ close to one:
(25) $\mathcal{L}_{unit}:=(\left\|h\right\|_{L^{2}(S)}-1)^{2}.$
@@
(26) $\mathcal{L}=\mathcal{L}_{Rayleigh}+\lambda_{ortho}\mathcal{L}_{ortho}+\lambda_{unit}\mathcal{L}_{unit}.$
@@
(27) $f,g\right>_{L^{2}(S)}\approx\frac{Area(S)}{N}\sum_{i=1}^{N}f(p_{i})g(p_{i}),$
@@
    Generate $M\gg N_{target}$ samples ($p_{i}$) from the 3D normal distribution $\mathcal{N}(\mathbf{\mu}=\mathbf{0},\mathbf{\Sigma}=\mathbf{I})$. Normalize, to unit norm, to produce a dense uniform distribution of points on the sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$.
@@
Following this process, the expected number of points will equal $N_{target}$ — a specified parameter — however, the actual number of points ($N$) may vary.
@@
Figure 7. Spectral Modes on Neural Surfaces. We propose how to compute the lowest few spectral modes of the neural Laplace Beltrami operator $\Delta_{S_{\theta}}$ via minimization of the Rayleigh quotient, without requiring any unnecessary discretization. Here we show the lowest few spectral modes of different SNSs. The lowest spectral mode is constant, and not included in this figure. Each of the spectral modes is represented by a dedicated MLP (see Equation 22).
Figure 8. Heat Flow (left) and Mean Curvature Flow (right). (Left) We can evolve a given scalar field, $f$, specified over the surface of an SNS, $S_{\theta}$, using the Partial Differential Equation (PDE) $f_{t}=\Delta_{S_{\theta}}f$ (the heat equation). We represent the evolving scalar field as a small MLP, whose weights are ‘finetuned’ for up to 100 epochs after every time step. Darker colors denote low values (cold) with lighter colors being high (hot). (Right) Taking advantage of the differentiable nature of our representation, we can also compute a mesh-free Mean Curvature Flow (MCF) in which the Mean Curvature and normals are updated at every iteration. We update the Spherical Neural Surface using a small number of finetuning steps, after every iteration of the flow. This formulation of MCF prevents singularities from forming, without any special handling to prevent them (Kazhdan et al., 2012).
@@
Figure 9. Laplace Beltrami on Neural Scalar Fields. We present two ways to compute Laplace Beltrami operators on any smooth scalar field, $f$, defined on a neural surface $S_{\theta}$ (top): first, using mean curvature estimates (top-middle) (Equation 15) and second, using the divergence of gradient definition (divgrad) (bottom-middle) (Equation 12). We represent scalar fields on surfaces using dedicated MLPs. (Bottom) For comparison, we also show results using the cotan LBO (with a lumped mass matrix) on a dense mesh, whose vertices are the image of the vertices in a dense icosphere mesh under the transformation $S_{\theta}$.

### v2

@@ 3.2 Computing Differential Quantities using SNS
$R=\begin{pmatrix}\cos\theta\cos\phi&-\sin\phi\\ \cos\theta\sin\phi&\cos\phi\\ -\sin\theta&0\\ \end{pmatrix}.$ (6)
The trigonometric quantities here are simple functions of $x$,$y$ and $z$.
This matrix is an orthonormal version of the usual spherical polar coordinates Jacobian. We avoid degeneracy at the poles by selectively re-labeling the $x$, $y$, and $z$ coordinates so that $\theta$ is never close to zero or $\pi$. (This is equivalent to choosing a different chart for the points within some radius of the poles.)
Now, from $\mathbf{J}^{\textit{local}}_{S}$, we can simply write the First Fundamental Form of the sphere-to-surface map as:
$\mathbf{I}_{S}=\begin{pmatrix}E&F\\ F&G\\ \end{pmatrix}=(\mathbf{J}^{\textit{local}}_{S})^{T}\,\mathbf{J}^{\textit{local}}_{S}.$ (7)
To compute the Second Fundamental Form, we first compute the second partial derivatives of the parametrization with respect to the local co-ordinates, and then find the dot products of the second derivatives with the normal, $\mathbf{n}=\mathbf{n}_{S_{\theta}}$. The second derivatives are:
$S_{uu}=\mathbf{J}(S_{u})S_{u},\quad S_{uv}=\mathbf{J}(S_{u})S_{v},\quad S_{vv}=\mathbf{J}(S_{v})S_{v}.$ (8)
We can now express the Second Fundamental Form as:
$\mathbf{II}_{S}=\begin{pmatrix}e&f\\ f&g\\ \end{pmatrix}=\begin{pmatrix}S_{uu}\cdot\mathbf{n}&S_{uv}\cdot\mathbf{n}\\ S_{uv}\cdot\mathbf{n}&S_{vv}\cdot\mathbf{n}\\ \end{pmatrix}.$ (9)
In terms of the Fundamental Forms, the Gauss curvature is
$K=\frac{eg-f^{2}}{EG-F^{2}},$ (10)
@@
$H=\frac{Eg-2Ff+Ge}{2(EG-F^{2})}.$ (11)
Furthermore, the solutions to the quadratic equation
$\text{det}(\mathbf{I}_{S}-\lambda\mathbf{II}_{S})=0$ give us explicit expressions for the principal curvatures, that we then use to find the principal curvature directions directly on the Spherical Neural Surface. In Figure 2, we compare our differential estimates against analytical computations for an analytic surface; in Figure 5 we also show curvature estimates and principal curvature directions on other SNSs overfitted to non-analytic surfaces (ground truth values are not available). In Section 6, we evaluate the quality of our estimates and also evaluate how the estimates change as we fit SNSs to different mesh resolutions as proxies to the same underlying (continuous) surface.
Figure 2: First and Second Fundamental Forms. We show an analytically defined surface, spike (bottom) and an SNS (top row) that has been overfitted to a mesh of the surface (with 40962 vertices). We show various differential quantities computed symbolically using Matlab (bottom) and computed on the SNS using our method (top). Our results align very closely with the analytically-computed results.
4 Laplace-Beltrami Operator
@@
4.1 Laplace-Beltrami Operator on Smooth Surfaces
The (Euclidean) Laplacian $\Delta$ of a scalar field $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is defined as the divergence of the gradient of the scalar field. Equivalently, $\Delta f$ is the sum of the unmixed second partial derivatives of $f$:
$\Delta f:=\nabla\cdot\nabla f=\sum_{i=1}^{n}\frac{\partial^{2}f}{\partial^{2}x_{i}}.$ (12)
@@ 4.1 Laplace-Beltrami Operator on Smooth Surfaces
$\Delta_{\Sigma}f:=\nabla_{\Sigma}\cdot\nabla_{\Sigma}f.$ (13)
The surface gradient of the scalar field $f$ (written $\nabla_{\Sigma}f$), and the surface divergence of a vector field ${\mathbf{F}}:\Sigma\rightarrow\mathbb{R}^{3}$ (written $\nabla_{\Sigma}\cdot\mathbf{F}$), can be computed by smoothly extending the fields to fields $\tilde{f}$ and $\tilde{\mathbf{F}}$ respectively, which are defined on a neighborhood of $\Sigma$ in $\mathbb{R}^{3}$.
The surface gradient is defined as the orthographic projection of the Euclidean gradient of $\tilde{f}$ into the local tangent plane:
$\nabla_{\Sigma}f=\nabla\tilde{f}-(\nabla\tilde{f}\cdot\mathbf{n})\mathbf{n},$ (14)
where $\mathbf{n}$ is the unit surface normal at the point. By choosing a suitable co-ordinate frame, one can show that this definition is independent of the choice of extension. The surface divergence of the vector field $\mathbf{F}$ is defined as
$\nabla_{\Sigma}\cdot\mathbf{F}=\nabla\cdot\tilde{\mathbf{F}}-\mathbf{n}^{T}{\mathbf{J}}(\tilde{\mathbf{F}})\mathbf{n},$ (15)
@@
By expanding out the definitions of surface gradient and surface divergence, we can derive an alternative formula for the surface Laplacian ($\Delta_{\Sigma}f$) in terms of the (Euclidean) Laplacian ($\Delta\tilde{f}$), the gradient ($\nabla\tilde{f}$) and the Hessian ($\mathbf{H}(\tilde{f})=\mathbf{J}(\nabla\tilde{f})^{T}$) as
$\Delta_{\Sigma}f=\Delta\tilde{f}-2H\nabla\tilde{f}\cdot\mathbf{n}-\mathbf{n}^{T}\mathbf{H}(\tilde{f})\mathbf{n}.$ (16)
The dependence on the surface is captured by the normal function $\mathbf{n}$, and the mean curvature $H$. (Refer to [Rei82, XZ03] for a derivation.) In the case when $\tilde{f}$ is a ‘normal extension’ — i.e., it is locally constant in the normal directions close to the surface — then the second and third terms disappear, so that $\Delta_{\Sigma}f$ is consistent with $\Delta\tilde{f}$.
If we substitute in the coordinate function $\mathbf{x}$ for $\tilde{f}$, then the first and third terms disappear, and we are left with the familiar equation
$\Delta_{\Sigma}\mathbf{x}=-2H\mathbf{n}.$ (17)
@@
As in section 3.2, these formulae are easily implemented for functions defined on an SNS. All that is required to implement equation 16 is that we have access to an extension $\tilde{f}$ and its first and second derivatives. Therefore it can be applied when ${f}$ is (the restriction to the surface of) a closed-form function on the embedding space (e.g. figures 10 and 11), or when $f$ and $\tilde{f}$ are defined by a network, allowing us to use autodiff (see section 4.2).
We have provided Equation 16) as an alternative formula to the standard definition (Equation 13) because where a field evolves on a domain that remains constant, it allows us to precompute the $\mathbf{H}$ and $\mathbf{n}$ and then only the Euclidean quantities ($\nabla\tilde{f}$, $\mathbf{H}(\tilde{f})$, and $\Delta\tilde{f}$) need to be updated. In Figure 8 (2nd and 3rd row) we demonstrate both methods and there is no visible difference.
4.2 Spectrum of the Laplace-Beltrami Operator
@@
Figure 3: Neural Spectral Basis against Analytical Solutions. The sphere is a genus-0 surface for which the eigenspaces of the Laplace-Beltrami operator are very well understood: the eigenspaces are the spaces spanned by linear combinations of the spherical harmonics for each frequency. This means that each of our optimized eigenfunctions should belong to one of these eigenspaces, and therefore, the Rayleigh quotient should match the energy level of the corresponding spherical harmonics. For evaluation, we compute the Rayleigh quotient using a large number of samples for Monte Carlo integration, since we cannot analytically compute the integral. For reference, we present the ground truth Rayleigh quotients, in green.
In the discrete mesh case, the eigenfunctions of the Laplace-Beltrami operator are computed by diagonalizing a sparse matrix [LZ10, BAB23] based on the input mesh. In the continuous case, however, we no longer have a finite-dimensional matrix representation for $\Delta_{\Sigma}$. Instead, we exploit a functional analysis result that describes the first $k$ eigenfunctions of $\Delta_{\Sigma}$ as the solution to a minimization problem, as described next.
@@
$Q_{\Sigma}(f):=\frac{\left\|\nabla_{\Sigma}f\right\|_{L^{2}(\Sigma)}^{2}}{\left\|f\right\|^{2}_{L^{2}(\Sigma)}}.$ (20)
@@
$\textit{such that}\quad\left<\Psi_{i},\Psi_{j}\right>_{L^{2}(\Sigma)}=0\quad\text{for all}\quad i\neq j.$ (21)
@@
$\Delta_{\Sigma}\Psi_{i}(x)=-Q_{\Sigma}(\Psi_{i})\Psi_{i}(x)\quad\forall\,x\in\Sigma.$ (22)
@@
Physically, the eigenfunctions of $\Delta_{\Sigma}$ represent the fundamental modes of vibration of the surface $\Sigma$. The Rayleigh Quotient of $\Psi_{0}$ is zero. As the Rayleigh quotient increases, the energy increases, and the eigenfunctions produce higher frequency patterns.
@@
Figure 4: Encoding Scalar Fields on Neural Surfaces. We encode scalar fields on the sphere (and corresponding extensions into $\mathbb{R}^{3}$) as MLPs $g_{\eta}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, with parameters $\eta$. Because $S_{\theta}$, considered as a parametrization, is bijective, then we can implicitly define any smooth scalar field $h$ on the surface in terms of one of the functions $g_{\eta}$ (restricted to $\mathbb{S}^{2}$). We visualize the one-to-one correspondence between scalar fields defined on the surface, and scalar fields defined on the sphere, by using vertex-colors to display the fields.
@@
Figure 5: First and Second Fundamental Forms. An SNS represents a smooth, seamless parametrization of a genus-0 surface, and we can compute the partial derivatives of this parametrization using automatic differentiation. We compute the normals and the First Fundamental Form of the parametrization from the first derivatives. Subsequently, computing the second derivatives enables us to construct the Second Fundamental Form of the parametrization. Hence, we compute Mean Curvature, Gaussian Curvature, and Principal Curvature Directions at any point without any unnecessary discretization error. To render scalar fields, we took samples at the 2.6 million vertices of an icosphere.
Figure 6: Accuracy of Differential Quantity Estimation. As shown by the colourmaps, the quality of the estimation of differential quantities (mean curvature, Gauss curvature, minimum curvature direction and normals) decreases gradually as mesh-resolution decreases. We show results for an SNS overfitted to analytical parametrisation, with 102400 sample points (top), and SNSs fitted to meshes with 40962, 10242 and 2562 vertices. (The errors are sampled at a much higher resolution, at 655362 points.) The level of error for the finest level mesh is close to the level of error for the ‘analytical’ SNS and we argue that this demonstrates the ‘interpolation’ capability of our method: it shows that fitting an SNS to an appropriately high-resolution mesh successfully interpolates the underlying surface (even for meshes where the analytic parametrisation is not available). We also show plots of the error on the various differential quantities, compared against three additional methods (when functions are available): Monge-fitting, NFGP and i3d. A minimum curvature direction function was not available for NFGP.
5 Spectral Analysis using SNS
@@
To find the continuous eigenmodes of the Laplace-Beltrami operator on an SNS, we require a continuous and differentiable representation for scalar functions defined on the surface. In our framework, we represent smooth scalar fields on the sphere $\mathbb{S}^{2}$ by MLPs, $g_{\eta}$, from $\mathbb{R}^{3}$ to $\mathbb{R}$, which are parametrized by $\eta$. We only ever evaluate the MLP $g_{\eta}$ on $\mathbb{S}^{2}\in\mathbb{R}^{3}$, but because the network’s domain is $\mathbb{R}^{3}$ then it defines an extension of the scalar field, and this allows us to compute the Euclidean gradient $\nabla g_{\eta}$. Then, if $S_{\theta}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ is an SNS, any smooth scalar field $h$ can be defined implicitly by the equation
$h\circ S_{\theta}=g_{\eta}.$ (23)
Applying the chain rule, we can compute the gradient of $h$:
$\nabla h=(\mathbf{J}(S_{\theta})^{T})^{-1}\nabla g_{\eta},$ (24)
@@
To optimize for a (smooth) scalar function $g_{\eta_{k}}$ so that $h$ approximates the $k$th non-constant eigenfunction of the Laplace-Beltrami operator, we use gradient descent to optimize the weights, $\eta_{k}$. Specifically, we use a combination of two loss terms:
$\mathcal{L}_{Rayleigh}:=Q_{S}(h)\quad\text{and}\quad\mathcal{L}_{ortho}:=\sum_{i=0}^{k-1}\left<h,h_{i}\right>_{L^{2}(S)}^{2},$ (25)
where we sequentially compute the eigenfunctions, and $h_{i}$ is the $i$th smallest eigenfunction. Recall that we represent each such eigenfunction using a dedicated MLP (except for $h_{0}$, which is the constant eigenfunction). The loss term $\mathcal{L}_{ortho}$ encourages the scalar field $h$ to be orthogonal to all previously-found eigenfunctions (including the constant eigenfunction). The quantity $Q_{S_{\theta}}(h)$ depends on the Jacobian of $S_{\theta}$, through $\nabla_{S_{\theta}}h$ and $\nabla h$. However, since the surface does not change during optimization and $\theta$ stays fixed, we precompute $(\mathbf{J}(S_{\theta})^{T})^{-1}$. Note that, if we were to optimize over the subspace of the unit-norm functions in $L^{2}(\Sigma)$, then we could replace the Rayleigh Quotient with the Dirichlet energy. However, it was a more natural design choice for us to parametrize the entire space $L^{2}(\Sigma)$ and optimize the Rayleigh Quotient, rather than strictly constrain the functions to have unit-norm and minimize the Dirichlet Energy.
Although functions $h$ are not required to have unit norm, we used a regularization term to prevent the $L^{2}$-norm of $h$ from either exploding or vanishing:
$\mathcal{L}_{reg}:=(\left\|h\right\|_{L^{2}(S)}-1)^{2}.$ (26)
@@
$\mathcal{L}=\mathcal{L}_{Rayleigh}+\lambda_{ortho}\mathcal{L}_{ortho}+\lambda_{reg}\mathcal{L}_{reg}.$ (27)
@@
$f,g\right>_{L^{2}(S)}\approx\frac{Area(S)}{N}\sum_{i=1}^{N}f(p_{i})g(p_{i}),$ (28)
@@
    Generate $M\gg N_{target}$ samples ($p_{i}$) from the 3D normal distribution $\mathcal{N}(\mathbf{\mu}=\mathbf{0},{\Sigma}=\mathbf{I})$. Normalize, to unit norm, to produce a dense uniform distribution of points on the sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$.
@@
Following this process, the expected number of sample points will equal $N_{target}$ — a specified parameter — however, the actual number of sample points ($N$) may vary.
@@
Figure 7: Spectral Modes on Neural Surfaces. We propose how to compute the lowest few spectral modes of the neural Laplace Beltrami operator $\Delta_{S_{\theta}}$ via minimization of the Rayleigh quotient, without requiring any unnecessary discretization. Here we show the lowest few spectral modes of different SNSs. The lowest spectral mode is constant, and not included in this figure. Each of the spectral modes is represented by a dedicated MLP (see Equation 23).
@@
Figure 8: Laplace Beltrami on Neural Scalar Fields. We present two ways to compute Laplace Beltrami operators on any smooth scalar field, $f$, defined on a neural surface $S_{\theta}$ (top): first, using mean curvature estimates (top-middle) (Equation 16) and second, using the divergence of gradient definition (divgrad) (bottom-middle) (Equation 13). We represent scalar fields on surfaces using dedicated MLPs. (Bottom) For comparison, we also show results using the cotan LBO (with a lumped mass matrix) on a dense mesh, whose vertices are the image of the vertices in a dense icosphere mesh under the transformation $S_{\theta}$.